rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#1533 — FS#5522 — boi-1-3k

Attached to Project— ADSL / SDSL
Incident
Backend / Core
CLOSED
100%
We have detected an optical problem:

TE0/2: Temperature high alarm; Operating value: 113.9 C, Threshold value: 100.0 C.

A technician is underway to replace the equipment. An interuption of a few minutes is expected during maintenance.


Date:  Tuesday, 14 June 2011, 19:34PM
Reason for closing:  Done
Comment by OVH - Tuesday, 14 June 2011, 19:32PM

The first optical bloc faces itself problems.

A thechnician is on the road to replace it.


Comment by OVH - Tuesday, 14 June 2011, 19:32PM

We are replacing the optical bloc.


Comment by OVH - Tuesday, 14 June 2011, 19:34PM

The second optical bloc is replaced. We renforce the monitoring
of this equipment for the next days in order to be sure of its
stability. The working threshold are currently normal.